OpenStudy (anonymous):

Can anyone answer what the quantity of x plus 4 over 3 = 2 is?

OpenStudy (anonymous):

bracket matters (xplus4)over 3 or xplus (4over3)

jigglypuff314 (jigglypuff314):

is your question \[\frac{ x+4 }{ 3 }=2\] or \[x + \frac{ 4 }{ 3 }=2\]

OpenStudy (anonymous):

It would be the first one

jigglypuff314 (jigglypuff314):

multiply 3 on both sides then subtract 4 on both sides

OpenStudy (anonymous):

Thank you!